OpenStudy (gabylovesyou):

Given the functions f(x) = 3x - 1 and g(x) = 3x + 4, which operation results in the smallest coefficient on the x term? Addition Subtraction Multiplication Two operations result in the same coefficient

OpenStudy (anonymous):

Hii i dmed the girl gaby on twitter but she said she doesn't have a openstudy? So your lying

OpenStudy (anonymous):

which one do you think it is?

OpenStudy (gabylovesyou):

@kyliejones which one did u pm ?

OpenStudy (gabylovesyou):

@kool_22 no idea

OpenStudy (anonymous):

The one in your bio :)

OpenStudy (anonymous):

Which one you think?

OpenStudy (gabylovesyou):

no u didnt lol cx

OpenStudy (anonymous):

Yeah i did lol your a fake

OpenStudy (gabylovesyou):

ok whatever.. @Kyliejones ;p i just checked my messages and got nun so ...

OpenStudy (anonymous):

Prove it.

OpenStudy (anonymous):

Are you chatting here? Or here to get help with your question?

OpenStudy (anonymous):

And I am water, so behave properly with me.. :P

OpenStudy (anonymous):

And the two functions and tell me, what will be the coefficient of \(x\) after adding?

OpenStudy (anonymous):

*Add.. very first word..

OpenStudy (gabylovesyou):

@eta tell her not me cx

OpenStudy (gabylovesyou):

and idk .-. @eta

OpenStudy (anonymous):

You cannot add the two functions or have not learnt yet?

OpenStudy (anonymous):

See, in adding, you just add the like terms of both functions: \[f(x) + g(x) = 3x - 1 \color{red}{+} \; 3x + 4\]

OpenStudy (anonymous):

So, what will you get when you add \(3x\) with \(3x\) ?

OpenStudy (gabylovesyou):

so 6x + 3

OpenStudy (anonymous):

Good, here, what is the coefficient of \(x\)?

OpenStudy (gabylovesyou):

6

OpenStudy (anonymous):

yep, remember it. Now subtract them..

OpenStudy (anonymous):

On subtracting, you will not have any \(x\) term, so this option is ruled out straightaway as you want to find smallest of coefficient of \(x\), and for that you definitely need \(x\)..

OpenStudy (anonymous):

Now, comes Multiplication.. \[f(x) \cdot g(x) = (3x - 1)(3x + 4) = ??\]

OpenStudy (gabylovesyou):

9x + 12x - 3x + 3

OpenStudy (anonymous):

Are you sure it is \(9x\)?

OpenStudy (anonymous):

Should not it be \(9x^2\) ?

OpenStudy (gabylovesyou):

ummm ohhhh ok

OpenStudy (anonymous):

What is the coefficient of \(x\) here?

OpenStudy (gabylovesyou):

9

OpenStudy (anonymous):

Good, so with which one you got smaller?

OpenStudy (gabylovesyou):

addition

OpenStudy (gabylovesyou):

@ganeshie8

OpenStudy (gabylovesyou):

idk how to divide them..

OpenStudy (anonymous):

Is there need to divide them?

OpenStudy (anonymous):

No choices you have with Divide operation.